OpenStudy (anonymous):

The tropical giant bamboo can grow 11.9 feet in 3 days. What is this rate of growth in inches per hour? Round your answer to the nearest hundredth, and show that the answer is reasonable.

OpenStudy (anonymous):

1 feet = 12 inches 11.9 feet=11.9*12=142.8 inches 1 day=24 hours 3 days = 24*3= 72 hours \[\frac{142.8}{72}\approx1.98333=1.98 inches per hour\]